A rectangle has a area of X^3-6X^2+15X-28 square feet. Its width is X-4 feet. What is its length?
Answer by edjones(8007)

A=L*w
L=A/w
(x^3-6x^+15x-28)/(x-4) Use long division or synthetic division.
=x^2-2x+7 Ans.
.
Ed